OpenStudy (anonymous):

A body hangs from a spring balance supported from the roof of an elevator. if the elevator has an upward acceleration of 3m/s^2 and the balance reads 50N.What is the true weight of the body...

OpenStudy (anonymous):

Using newtons seconds law F=m*a you should be able to solve this

OpenStudy (anonymous):

Note: Weight=Force

OpenStudy (anonymous):

W=M*G is the same as F=M*A

OpenStudy (anonymous):

U mean Mass is 50/3? @Brent0423

OpenStudy (anonymous):

@experimentX ANY IDEAS?

OpenStudy (experimentx):

f = m(a + g)

OpenStudy (anonymous):

PLS ANY EXPLAINATION WHY IS DAT BECAUSE @experimentX THE ANS U SUGGEST IS NOT IN OPTIONS

OpenStudy (anonymous):

THE OPTIONS ARE 50, 28.3, 38.3 AND 39.8

OpenStudy (experimentx):

50 = m(3 + 9.8) find the mass m true weight is m x 9.8

OpenStudy (experimentx):

or 50/12.8 x 9.8 = 38.2813

OpenStudy (anonymous):

SO ANS IS 38.3

OpenStudy (experimentx):

yup

OpenStudy (anonymous):

THANKS @experimentX PLS EXPLAIN WHY WE HAVE TO ADD (A + G)

OpenStudy (experimentx):

 this will increase relative acceleration which will force